# Understanding Sampler Stochasticity in Training Diffusion Models for RLHF

### v1

@@ Appendix D Proof for Section 4
Since $\alpha(t)$ satisfies the $\kappa$-strict negativity by assumption, we may apply equation 18:
$u(T)$ $\leq\int_{0}^{T}e^{-\kappa(T-s)}\cdot\beta(s)ds$
@@
Finally, $W_{2}$ distance is given by taking square root of $L^{2}$ distance between $Y^{ODE}$ and $Y^{SDE}$.
Appendix E Hyperparameters
@@
E.1 DDPO Experiments
We follow the setup of Black et al. (2024), using denoising step $T=50$ and guidance weight $w=5.0$ throughout all experiments. We also use the AdamW optimizer Loshchilov & Hutter (2019) with default weight decay 1e-4 and optimal learning rates for different reward functions. Reward gaps under four reward functions are shown in Figure 3 in Section 5 and Appendix G.
@@ E.1 DDPO Experiments
Animal Prompts dataset consists of 398 animal labels extracted from ImageNet-1k class labels Deng et al. (2009), often with comma-separated synonyms and scientific names.
@@
E.2 MixGRPO Experiments
We follow the setup of Li et al. (2025), letting reward model be ”multi_\_reward” with equal weights. We set $T=15$ as the denoising steps, AdamW optimizer with learning rate 1e-5 and weight decay 1e-4. For GRPO, the generation group size is 12 and clip range is 1e-4. We perform 12 gradient updates per iteration.
Appendix F LLM Usage
@@
Appendix G More Experiment Results
G.1 DDPO Images on different training steps
@@
G.2 DDPO Images under ImageReward
@@
G.3 DDPO Images under PickScore
@@
G.4 DDPO Reward Gaps for other rewards
G.4.1 ImageReward

### v2

@@ D.2 Proof for strongly log-concave distributions
Theorem D.1.
[…]
Proof.
With the additional log-concavity assumption on the score function,
$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}}))\,\big\rangle\leq-\kappa\cdot u(t).$
In this case, the coefficient $\alpha$ becomes contractive as we may take $\delta=\frac{\kappa-1}{2\kappa}$:
@@
For arbitrary time horizon $T$, equation 18 gives:
$u(T)$ $\leq\frac{(1-e^{-\alpha T})\cdot\beta}{-\alpha}$
[…]
Remark.
Assumption 4.2 is valid for a unimodal terminal distribution, in which a strongly log-concave coefficient $\kappa(T)$ satisfies
$\mathbb{E}\big\langle\,Y_{T}^{\text{ODE}}-Y_{T}^{\text{SDE}},s_{\theta}(T,Y_{T}^{\text{ODE}})-s_{\theta}(T,Y_{T}^{\text{SDE}})\,\big\rangle\leq-\kappa(T)\cdot u(T).$
In intermediate time steps,
$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}}))\,\big\rangle\leq-\kappa(t)\cdot u(t),$
in which (TZ24tut) bounds
$\kappa(t)$ $\geq\frac{\kappa(T)}{\exp\left(-\int_{0}^{T-t}f(s)\,ds\right)+\kappa(T)\cdot\int_{0}^{T-t}\exp\left(-\int_{s}^{T-t}f(\tau)d\tau\right)ds}.$
Since $\kappa(t)$ is strictly positive on $[0,T]$, a global coefficient $\kappa_{\text{global}}=\inf_{t}\kappa(t)$ exists.
D.3 Discretization Error Analysis
Let $N$ be the number of time steps and $h=T/N$ be the step size, we consider the DDPM sampling scheme with Euler–Maruyama discretization. Our Assumption 4.1 are equivalent to Assumptions 6.1 and 6.2 ($L$-Lipschitz score and finite second moment) in (gaussianmixture), which develops from (chen202). For $h\lesssim 1/L$,
$\operatorname{TV}(q_{T}^{h},p_{0})\;\lesssim\;\underbrace{\sqrt{\mathrm{KL}(p_{0}\|\mathcal{N}(0,I))}e^{-T}}_{\text{forward convergence}}\;+\;\underbrace{(L\sqrt{dh}+Lm_{2}h)\sqrt{T}}_{\text{discretization error}}\;+\;\underbrace{\varepsilon_{0}\sqrt{T}}_{\text{score error}},$
where $q_{T}^{h}$ is the law of the discrete sampler at time $T$, $d$ is the ambient space dimension, and $m_{2}^{2}=\mathbb{E}_{p_{0}}\|X\|^{2}$. In particular, the discretization contribution to the total variation distance is $O(h^{1/2})$ as $h\to 0$ for fixed $T,d,L,m_{2}$.
Denote $Y_{T}^{\text{ODE},h}$ and $Y_{T}^{\text{SDE},h}$ as the ODE/SDE fine-tuned inference under step size $h$, by triangular inequality,
$W_{2}(Y_{T}^{\text{ODE},h},Y_{T}^{\text{SDE},h})\leq W_{2}(Y_{T}^{\text{ODE}},Y_{T}^{\text{SDE}})+\underbrace{W_{2}(Y_{T}^{\text{ODE},h},Y_{T}^{\text{ODE}})+W_{2}(Y_{T}^{\text{SDE}},Y_{T}^{\text{SDE},h})}_{\epsilon_{d}(h)}$
Since $\sup_{t}\mathbb{E}\|X_{t}\|^{2}\leq\infty$, a standard coupling method between $W_{2}$ and $\operatorname{TV}$ distance gives.
$W_{2}(Y_{T}^{\theta},Y_{T}^{\theta,h})\;\;\lesssim\;\;\operatorname{TV}(Y_{T}^{\theta},Y_{T}^{\theta,h})^{1/2}.$
Therefore, the contribution of time discretization to the $W_{2}$ error is bounded by the additive term
$\epsilon_{d}(h):=C\,\bigl(L\sqrt{dh}+Lm_{2}h\bigr)^{1/2},$
for some constant $C$ depending only on $T,d$. In the continuous time limit, $\epsilon_{d}(h)\to 0.$
Appendix E Discussion on DDIM and gDDIM under high stochasticity
When $1.0<\eta<+\infty$, the stochasticity level of gDDIM is always upper-bounded by
$\sigma_{t}^{\text{gDDIM}}(\eta)=(1-\alpha_{t-\Delta t})\left(1-\left(\frac{1-\alpha_{t-\Delta t}}{1-\alpha_{t}}\right)^{\eta^{2}}\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\right)^{\eta^{2}}\right)\leq 1-\alpha_{t-\Delta t}<+\infty;$
whereas the linear interpolation of DDIM gives an unbounded
$\sigma_{t}^{{\text{DDIM}}}(\eta)=\eta\sqrt{\frac{1-\alpha_{t-\Delta t}}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}.$
In order to ensure the well-posedness of $\sqrt{(1-\alpha_{t-\Delta t}-\sigma_{t}^{2})(1-\alpha_{t})}$ in equation 4, the conventional discretization requires
$\sigma_{t}^{\text{DDIM}}\leq 1-\alpha_{t-\Delta t}.$
We may attempt to bypass the restriction by regulating
$\sigma_{t}^{\text{linear DDIM}}(\eta)=\min\{1-\alpha_{t-\Delta t},\eta_{t}^{\text{DDIM}}(\eta)\}.$
However, this changes the terminal marginals in equation 3. Therefore, the classical DDIM interpolation cannot simultaneously support arbitrary $\eta>1.0$ and preserve the terminal marginal, whereas gDDIM does.
Appendix F Hyperparameters
@@
F.1 DDPO Experiments
We follow the setup of ddpo, using denoising step $T=50$ and guidance weight $w=5.0$ throughout all experiments. We also use the AdamW optimizer adamw with default weight decay 1e-4 and optimal learning rates for different reward functions. Reward gaps under four reward functions are shown in Figure 3 in Section 5 and Appendix H.
@@ F.1 DDPO Experiments
Animal Prompts dataset consists of 398 animal labels extracted from ImageNet-1k class labels imagenet, often with comma-separated synonyms and scientific names.
@@
F.2 MixGRPO Experiments
We follow the setup of MixGRPO, letting reward model be ”multi_\_reward” with equal weights. We set $T=15$ as the denoising steps, AdamW optimizer with learning rate 1e-5 and weight decay 1e-4. For GRPO, the generation group size is 12 and clip range is 1e-4. We perform 12 gradient updates per iteration.
Appendix G LLM Usage
@@
Appendix H More Experiment Results
H.1 DDPO Images on different training steps
@@
H.2 DDPO Images under ImageReward
@@
H.3 DDPO Images under PickScore
@@
H.4 DDPO Images under more comprehensive prompts
SDE Refer to caption Refer to caption Refer to caption Refer to caption
ODE Refer to caption Refer to caption Refer to caption Refer to caption
H.5 DDPO Reward Gaps for other rewards
H.5.1 ImageReward